OpenStudy (anonymous):

I need help!! can somebody solve for t using 980t^2/4(3.14)^2

OpenStudy (campbell_st):

well you don't have an equation so you can't solve it. It has to be equal to something....

OpenStudy (anonymous):

it is equal to 35

OpenStudy (campbell_st):

ok so its \[\frac{980t^2}{4(3.14)^2} = 35 \] start by multiplying both sides of the equation by \[4(3.14)^2 \] this will allow you to get rid of the denominator when you have an answer... divide both sides by 980 then you are left with \[t^2 = \frac{4(3.14)^2 \times 35}{980}\] last step is to take the square root of both sides.

OpenStudy (anonymous):

is it 31.55 im really bad at math